### v1

# Spin excitations of the Shastry-Sutherland model – altermagnetism and proximate deconfined quantum criticality

Abstract
Symmetry plays a crucial role in condensed matter physics. In quantum magnetism, it dictates a number of exotic phenomena, including deconfined quantum criticality and altermagnetism. Here, by studying the spin excitations of the $S=1/2$ antiferromagnetic Shastry-Sutherland model, we show that the Néel antiferromagnetic state in this model is an altermagnet featuring a non-relativistic splitting between two chiral magnon bands. Moreover, we identify a Higgs mode in the longitudinal excitation channel, whose gap softens when approaching the antiferromagnetic to plaquette valence bond solid transition, implying the appearance of nearly deconfined excitations. However, the splitting between the two magnon bands (Goldstone modes) remains finite at the transition. These results indicate that the transition is weakly first-order and proximate to a putative deconfined quantum critical point. We find that the altermagnetism provides a sensitive means to probe the deconfined quantum criticality.
Introduction.—Symmetry is essential in physics. It provides an organizational principle to classify complex states of matter, and based on this, many exotic phenomena are understood TI_Review ; Senthil_AnnRevCondMattPhys_2015 ; Witten_NP_2018 . One prominent example is the recently proposed altermagnet, a collinear antiferromagnetic (AFM) state in which the two sublattices with oppositely polarized spins are connected by the rotational symmetry instead of inversion or translation Smejkal_PRX_2022A ; Smejkal_PRX_2022B . The distinct symmetry, which is organized by the spin space group theory Yang_arXiv:2105.12738 , causes a non-relativistic spin-dependent splitting of energy bands and may give rise to some exotic effects, such as anomalous Hall effect Smejkal_PRX_2022A ; Smejkal_NatRevMater7_2022 . Up to now, most studies focus on itinerant systems Gonzalez_Betancourt_PRL_2023 ; Hariki_PRL_2024 ; Ferrari_arXiv:2408.00841 . For insulating altermagnets, though a symmetry dictated splitting in magnon bands has been predicted Smejkal_PRL131_256703_2023 , existing studies are mostly limited to several model systems Ma_PRB110_064426_2024 ; Weissenhofer_PRB110_094427_2024 ; Yao_arXiv_2024 ; Zhaojize_arxiv .
Symmetry can also dictate phase transitions. In the Landau-Ginzburg-Wilson paradigm, a system transitions from a disordered state to an ordered one by spontaneously breaking a symmetry. With quantum fluctuations the transition can be continuous and occur at zero temperature, giving rise to intriguing quantum critical phenomena sachdev_2011 ; Coldea_2010 ; Senthil2004 ; Cui2023DQCP ; Cui2019SCVO ; Lee2019 ; E8_2021 ; Xu2022 ; agrawal2020quantum ; fey2019quantum ; zhang2011exploring . For a transition between two ordered phases, the Landau paradigm predicts a first-order transition. However, this scenario has recently been challenged. It was proposed that the transition between a valence bond solid (VBS) and an antiferromagnetic (AFM) phase can be continuous, e.g., via a deconfined quantum critical point (DQCP) Senthil2004 . At the DQCP, the establishment of one order is accompanied by the destruction of the other order via proliferation of topological excitations, giving rise to emergent deconfined fractionalized excitations as well as enhanced continuous symmetry among the seemingly unrelated two types of order parameters. This scenario has motivated extensive studies on DQCP Sandvik2007 ; Meng_EasyplaneDQCP ; Shao_Science ; Senthil_Vojta_Sachdev:2004 ; Liu_Vojta:2022 ; Guo_PRL_2024 . However, whether a DQCP can be realized in quantum spin models is still an open question.
Tuning the frustration is a promising way to induce a VBS-to-AFM transition, and a prominent example is given by the Shastry-Sutherland (SS) model Shastry1981 . The model is defined on the SS lattice as illustrated in Fig. 1(a). It contains orthogonal spin dimers (with coupling $J^{\prime}$) connected by frustrated nearest neighbor (n.n.) bonds (with coupling $J$). With increasing $J/J^{\prime}$, the ground state experiences a series of quantum phase transitions from a dimer singlet phase to a plaquette VBS at $J/J^{\prime}\approx 0.68$, then to a Néel AFM state for $J/J^{\prime}\sim 0.8$ (see Fig. 1(b)).Koga2000 ; Chung2001 ; Pixley2014 ; Corboz2013 ; Boos2019 ; Lee2019 ; yang2021quantum ; Sengupta_PRL:2013 ; Wang_PRL_2023 . Although the dimer-to-plaquette transition is clearly strongly first-order, the nature of the plaquette-to-AFM transition remains elusive: Several tensor network calculations Corboz2013 ; Xi2023 showed it to be weakly first-order, while some other study yang2021quantum found an intervening gapless quantum spin liquid in between the plaquette and AFM phases. Interestingly, one DMRG work Lee2019 suggested the transition is via a DQCP with an emergent $O(4)$ symmetry, and quantum scaling behavior near this point has been revealed by a finite-size tensor network analysis WenyuanLiu_PRL_2024 . One prominent signature of a DQCP is the emergence of deconfined fractionalized spin excitations. Although spin excitations in both plaquette and AFM states have been extensively studied Mila_BOT ; Pinaki_PRB92_094440_2015 ; WangZ2018 ; Wang2023 , whether deconfined excitations emerge at the transition is still unclear.
The phase diagram of the SS model well describes the evolution of low-temperature phases of the quasi-2D antiferromagnet $\mathrm{SrCu}_{2}(\mathrm{BO}_{3})_{2}$ under pressure tuning Kageyama1999 ; PhysRevLett.82.3701 ; Koga2000 ; Waki2007 ; haravifard2016crystallization ; Zayed2017 ; Bettler2020 ; Guo2020 ; Jimenez2021 , giving the hope to realize a DQCP in this material. Remarkably, a recent NMR study in this compound Cui2023DQCP reported an unusually large anomalous dimension $\eta$ at a field-induced plaquette-to-AFM transition under pressure, implying behavior in proximity to a DQCP. However, direct spectral evidence on the deconfined excitations at zero field is absent.
In this Letter, we study spin excitations of the SS model by using the state-of-the-art tensor network method for infinite lattice systems. This prevents from the severe finite-size effect in usual numerical calculations and offers a new perspective on the nature of emergent DQCP phenomena. We show that the altermagnetic nature of the AFM state causes a finite splitting with opposite chirality in the magnon bands along the $\Gamma$ to M direction of the Brillouin zone. Besides the transverse excitations, we also identify a Higgs mode in the longitudinal excitation channel, whose energy decreases with reducing $J/J^{\prime}$ ratio. Approaching the plaquette-to-AFM transition, the gaps of the Higgs mode and the triplet excitations in the plaquette phase both rapidly drop close to zero, implying the emergence of nearly deconfined excitations. However, the splitting of the magnon bands remains finite at the transition. These results suggest the transition is weakly first-order, in proximity to a DQCP with an emergent O(4) symmetry. In general, we show that the split magnon bands caused by the altermagnetism serve as a sensitive probe for a deconfined quantum phase transition.
Figure 1: (a): Sketch of the SS lattice and the corresponding altermagnetism. The two sublattices of the Néel state are connected by a $C_{4}$ rotation of the lattice about the center of the empty plaquette (without the dimer bond). A time-reversal symmetry then recovers the Néel order. (b): Sketched phase diagram of the SS model with $J/J^{\prime}$. The $J/J^{\prime}$ values at the two transitions are taken from Ref. Xi2023 . (c): LSW dispersion in the AFM state of the SS model showing the chiral magnon bands with a non-zero splitting along the M-$\Gamma$-M direction of the Brillouin zone. $\mathcal{L}$($\mathcal{R}$) refers to the left(right)-handed chiral magnon bands.
Model and method.—The Hamiltonian of the SS model reads as
$\mathcal{H}=J\sum_{\langle i,j\rangle}\mathbf{S}_{i}\cdot\mathbf{S}_{j}+J^{\prime}\sum_{\langle\langle i,j\rangle\rangle}\mathbf{S}_{i}\cdot\mathbf{S}_{j},$ (1)
[…]
Figure 2: (a): Spin DSF in the AFM phase of the SS model calculated by the LSW theory. The inset shows high symmetric directions of the Brillouin zone. (b): Total spin DSF ($\mathcal{S}^{xx+yy+zz}$) in the AFM phase calculated by the iPEPS method with $D=3$ at $J/J^{\prime}=1.1$. (c): The transverse component ($\mathcal{S}^{xx+yy}$) of the results in (b). (d): The longitudinal component ($\mathcal{S}^{zz}$) of the results in (b). The dashed lines in (b) and (d) label the Higgs mode.
Altermagnetism and chiral magnons in the SS model.—The space group of the SS lattice is non-symmophic and the two sublattices of the Néel AFM state are connected by neither translational nor inversion, but a $C_{4}$ rotation about the center of the empty plaquette. The AFM order then recovers by further applying a time-reversal $\mathcal{T}$ symmetry to the spins (see Fig. 1(a)). This indicates that the Néel AFM state on the SS lattice is an altermagnet with a $d$-wave symmetry. The spin space group symmetry dictates a non-zero splitting with opposite chirality of the magnon bands along the diagonal directions of the Brillouin zone. This is confirmed by the LSW calculation shown in Fig. 1(c). Along the M-$\Gamma$-M direction, the acoustic (Goldstone) modes split into two branches with left- and right-handed chiralities, respectively. The splitting between the two chiral modes along $\Gamma$-M and $\Gamma$-M directions has opposite sign, and its magnitude reaches a maximum at the mid-point $\Sigma$ (and $\Sigma^{\prime}$). On the other hand, the two modes are degenerate along the $\Gamma$-X (and X-M) directions (Fig. 2(a)). These properties verify that the altermagnet has a $d$-wave nature. Note that the optical modes also split, but one band is completely flat along the M-$\Gamma$-M due to frustration and has zero spectral weight.
Spin excitation spectra.—We implement tensor network calculation to go beyond the LSW results on the spin excitation spectra. The calculated spectrum at $J/J^{\prime}=1.1$ is shown in Fig. 2(b). It exhibits richer structures than the LSW results in Fig. 2(a). To understand these structures, we calculate the transverse and longitudinal components of the DSF, $\mathcal{S}^{xx+yy}(\mathbf{q},\omega)$ and $\mathcal{S}^{zz}(\mathbf{q},\omega)$, and show them in Fig. 2(c) and (d), respectively. The transverse DSF resembles the LSW spectrum with three dispersive magnon bands. The finite gap at M (and $\Gamma$) of the Goldstone modes is an effect of finite bond dimension $D$ of the iPEPS ansatz, and vanishes in the large-$D$ limit (Fig.S3 of SM SM ). One prominent feature is the splitting of Goldstone modes along the $\Gamma$-M direction. The opposite sign of $\Delta\mathcal{S}=\mathcal{S}^{+-}(\mathbf{q},\omega)-\mathcal{S}^{-+}(\mathbf{q},\omega)$ for the two modes shown in Fig.S4 justifies their opposite chirality, as expected for an altermagnet.
The spectrum of longitudinal modes, shown in Fig. 2(d), also consist of multiple bands, owing to the complex SS lattice structure (see discussion below). We can identify a mode (labeled by the dashed line) that develops a strong resonance at the M point. The energy of this mode decreases with reducing $J/J^{\prime}$ (see Fig.S2 of SM SM ). We then attribute it to the Higgs mode which reflects the amplitude fluctuations of the AFM order parameter. In a two-dimensional AFM state, the Higgs mode is usually hardly visible because the gapless Goldstone modes give rise to a divergent longitudinal susceptibility and the Higgs mode can decay into a pair of Goldstone modes Poldolsky_PRB84_174522_2011 . Here it shows up in the tensor network calculation for the following reasons: The gapped Goldstone modes at finite $D$ remove the divergence, and the single-mode approximation adopted suppresses the two-magnon continuum but preserves the two-magnon bound state, which is just the Higgs mode. Just above the Higgs mode, we find another longitudinal mode that is quasi-degenerate to the Higgs one at M point but develops a sizable separation at $\Sigma$ point. Similar behavior was reported in a related model Zhaojize_arxiv , and its origin needs further study. Besides the collective excitation modes discussed above, the spectrum also contains a continuum, as shown in the broad peak above about $2J^{\prime}$ in the density of states in Fig. S5 of SM SM . Interestingly, the peak moves to lower energy when reducing $J/J^{\prime}$ toward the plaquette-to-AFM transition, in a way similar to the behavior of the Higgs mode.
Figure 3: (a): Spin DSF in the plaquette phase calculated by the iPEPS method with $D=3$ at $J/J^{\prime}=0.7$. The red dashed line labels the dispersion of the triplet excitations in the bond-operator theory. (b): $J/J^{\prime}$ dependence of extracted gaps at the M point of the Brillouin zone for several low-energy excitation modes in the plaquette phase, including, from bottom to top, a triplet, a singlet, and a 7-fold multiplet. The results are obtained by the iPEPS method with $D=4$. The dashed line shows the gap of the triplet excitations in the bond-operator theory.
For a comprehensive understanding, we also study the excitations in the plaquette phase, and a typical spectrum at $J/J^{\prime}=0.7$ is shown in Fig. 3(a). The lowest-energy excitations are found to be a triplet mode, whose dispersion can be described by the bond-operator theory Mila_BOT ; SM . Fig. 3(b) illustrates excitation gaps at the M point of all identified low-energy modes, which, from bottom to top, include the triplet mode, one singlet mode, and a 7-fold multiplet. Note that the minor splitting probably comes from numerical accuracy and can be eliminated by enforcing the spin SU(2) symmetry in the calculation. The gap of the triplet mode decreases with increasing $J/J^{\prime}$ toward the plaquette-to-AFM transition and the gap size is comparable with that from the bond-operator theory. The energy of the singlet mode is almost independent of $J/J^{\prime}$, and that of the 7-fold multiplet slightly increases with $J/J^{\prime}$. We calculate the quantum numbers of the 7-fold multiplet and find that this multiplet consists of two triplets and one singlet, which are the eigenstates of a single plaquette(see Fig. S6 of SM SM ). The inter-plaquette interaction then couples multiplets in each plaquette into bands that can be understood by a generalized bond-operator theory Pinaki_PRB92_094440_2015 . Note that each triplet excitation will further split to transverse and longitudinal modes once the magnetic order breaks the spin SU(2) symmetry, accounting for the complex multiband structure shown in Fig. 2 in the AFM phase. The singlet excitation cannot be described within a single plaquette picture and is interpreted as an inter-plaquette bound state. Above these low-energy excitations, the spectrum becomes continuous above about $J^{\prime}$ (Fig. S5 of SM SM ), similar to the case of the AFM phase.
Figure 4: (a): Gaps at the M point of the triplet mode (TM), Higgs mode (HM), and Goldstone modes (GM) with $J/J^{\prime}$ near the plaquette-to-AFM transition, obtained in the iPEPS calculation with $D=4$. (b): Finite-$D$ analysis on the crossing points of the three types of excitation modes in (a). In the large-$D$ limit, all crossing points are close to the plaquette-to-AFM transition at $(J/J^{\prime})_{c}\approx 0.79$, suggesting the transition is proximate to a (putative) DQCP. (c): Extrapolation of the energy of the crossing point between the Higgs mode and the triplet mode with $1/D$. The small energy in the large-$D$ limit indicates existence of nearly deconfined excitations. (d): Evolution of spin-wave velocities $c$ and the anisotropic parameter $w$ of the two Goldstone modes with $J/J^{\prime}$ in the AFM phase from the iPEPS calculation. Approaching the plaquette-to-AFM transition (indicated by the vertical dashed line), $w$ remains to be finite, indicating non-degenerate Goldstone modes at the transition point.
Probing the DQCP.—With excitations studied in both plaquette and AFM phases, we now understand the nature of the plaquette-to-AFM transition. One signature of DQCP is that at the transition, the triplet mode and Higgs mode should both become gapless and degenerate to the Goldstone modes. In Fig. 4(a) we plot the extracted gaps of the triplet mode, Higgs mode, and Goldstone modes at the M point with bond dimension $D=4$. The gaps for other $D$ values are presented in Fig. S7 of the SM SM . The gap of the Goldstone modes is almost independent of $J/J^{\prime}$, indicating that the gap is caused by finite $D$. On the other hand, gaps of both triplet mode and Higgs mode drop rapidly and are close to the gap of the Goldstone modes when approaching the transition point $(J/J^{\prime})_{c}\approx 0.79$.Xi2023 The crossing points of these modes are all close to $(J/J^{\prime})_{c}$ when extrapolating to the large-$D$ limit, as shown in Fig. 4(b). These behaviors imply proximity to a continuous quantum phase transition with nearly deconfined excitations. However, detailed analysis shows the crossing point of Higgs mode and Goldstone modes is at $J/J^{\prime}<(J/J^{\prime})_{c}$ at $D\rightarrow\infty$, while the one of triplet mode and Goldstone modes is at $J/J^{\prime}>(J/J^{\prime})_{c}$. This suggests that both Higgs mode and triplet mode are still gapped when extrapolating to the transition point. Indeed, as shown in Fig. 4(c), the energy of the crossing point between Higgs mode and triplet mode is about $0.05J^{\prime}$, a small gap close to zero.
Interestingly, we find the nature of the plaquette-to-AFM transition can be detected by the splitting of the two Goldstone modes. The DQCP theory for the SS model proposed the emergence of an O(4) symmetry between the AFM and plaquette order parameters Lee2019 . According to this theory, the Higgs mode and the two Goldstone modes would be degenerate and become the lower edge of the deconfined spinon continuum $\omega_{\text{min}}(\mathbf{q})\sim J|\mathbf{q}|$ at the DQCP. However, because the Néel AFM state of the SS model is an altermagnet, splitting between the two chiral Goldstone modes should be present when the AFM order parameter is finite. The splitting at the transition point vanishes only if the transition is continuous, via a DQCP. We fit the dispersions of the two Goldstone modes along the $\Gamma$-M direction by $\omega_{\pm}(\mathbf{q})=\sqrt{\Delta^{2}+c^{2}q^{2}}\pm wq_{x}q_{y}$, where $c$ refers to the spin-wave velocity in the gapless ($\Delta\rightarrow 0$) limit. The $w$ term accounts for the split Goldstone modes in the altermagnet with a $d$-wave symmetry where $w$ is an anisotropic parameter. The fitted $c$ and $w$ parameters are shown in Fig. 4(d). Approaching the transition, both of them are finite. This indicates that well-defined Goldstone modes with a non-zero splitting persist at the transition point. Together with the small gap of Higgs mode approaching the transition, the numerical results suggest the plaquette-to-AFM transition is weakly first-order and proximate to a DQCP.
Discussions and conclusions.—To understand the numerical results, note that the splitting of the Goldstone modes is caused by the $d$-wave anisotropic $w$ term in the $q^{2}$ order of the magnon dispersions. The Goldstone modes are still degenerate up to the linear-in-$q$ term. Therefore, one argues that the $w$ term, though a symmetry-breaking perturbation, is irrelevant in the low-energy effective theory, which predicts a DQCP with an emergent O(4) symmetry. However, given that the O(4) DQCP is obtained by deforming an SO(5) conformal field theory (CFT) along a particular direction Lee2019 and the SO(5) CFT could be non-unitary ma2020theory ; nahum2020note ; zhou2024so , realization of an O(4) DQCP on a lattice model would be delicate. Most likely, perturbations reflecting the exact lattice symmetry would eventually make the DQCP proximate.
This proximate DQCP scenario is supported by the appearance of rich low-lying modes on both sides when approaching the transition. These include the softening of the triplet mode in the plaquette phase and the Higgs mode in the AFM phase. In addition, we also observe a singlet excitation in the plaquette phase and quasi-degenerate longitudinal modes in the AFM phase. The appearance of more than three low-energy excitations on either side of the transition suggests that it goes beyond the conventional O(3) Wilson-Fisher universality or an O(4) symmetry-enhanced first-order transition, but is instead proximate to a putative O(4) DQCP, which supports both soft $S=1$ and $S=0$ modes. Moreover, our scenario naturally applies to a large class of models including the SS model, the Heisenberg model on the checkerboard lattice, the checkerboard $J$-$Q$ model, and their variants, all of which hold a $d$-wave altermagnetic phase dictated by the $C_{4}\mathcal{T}$ symmetry. The proximity to a putative DQCP provides a unified explanation for the observed quantum scaling behaviors at the plaquette-to-AFM transition in both the SS model WenyuanLiu_PRL_2024 and the checkerboard $J$-$Q$ model Zhao_NP_2018 ; ChengchenLi_JPCM_2024 . It would be interesting to understand how the spin frustration influences the transition by systematically studying the excitation spectra of these models.
We have examined the lattice structure of the SS model material SrCu2(BO3)2, and find that the structure indeed supports an altermagnet even when the non-magnetic ions are included. Our proximate DQCP scenario then applies: The transition is weakly first-order and proximate to a putative DQCP. Indeed this has been observed in a recent NMR experiment under finite magnetic field Cui2023DQCP and a study on the specific heat at zero field LilingSun_arXiv_2024 .
Besides the DQCP, another unsettled issue is the nature of the plaquette state. Recent inelastic neutron scattering (INS) Zayed2017 and NMR Cui2023DQCP measurements suggest a full plaquette phase where the singlet is located on the plaquette with a $J^{\prime}$ dimer. But most numerical results (including ours) on the SS model found the ground state is an empty plaquette (without the $J^{\prime}$ dimer). Interestingly, for the plaquette state at $J/J^{\prime}=0.7$, if taking $J^{\prime}\sim 4$ meV estimated from the experiment, the calculated spin gaps of the lowest triplet mode and the 7-fold degenerate mode (including two degenerate triplets) are about 1 meV and 2 meV, which are consistent with the gap values of the lowest two excitations observed in the INS experiment Zayed2017 . It would be important to further compare the theoretical and experimental results on the dispersions and evolution of the spin gaps with pressure.
To experimentally probe the altermagnetism, one may look for the splitting of the two magnon bands (at the order of $J^{\prime}$ near the $\Sigma$ point) in the spectrum of INS. The chiral magnons can be detected by resonant inelastic X-ray scattering (RIXS). Moreover, the transverse spin current induced by the chiral magnons will cause sizable spin Nernst effect in the AFM state Smejkal_PRX_2022B ; Cui_PRB_2023 when applying a temperature gradient along the non-splitting direction. Given that the Dzyaloshinsky-Moriya interaction is much smaller than $J$ and $J^{\prime}$ in SrCu2(BO3)2, the leading contribution should be attributed to the altermagnetism. Given that the splitting of the magnon bands in the altermagnetic phase is directly related to the O(4) symmetry breaking in the magnetic sector, the proposed experiments above, provide viable ways to detect the proximate deconfined quantum criticality.
In conclusion, we study the spin excitations of the SS model by using advanced tensor network method. We show that the antiferromagnetic state of the model is an altermagnet supporting two split magnons (Goldstone modes) with opposite chirality. Besides the Goldstone modes, we also identify a Higgs mode in the longitudinal channel. The spin excitations in the plaquette singlet phase contain several triplet modes. The dispersion of the lowest-energy triplet mode can be described by the bond-operator theory. When the system approaches the transition, rich excitations appear in low energies, and the gaps of both the Higgs and the triplet excitations reduce rapidly close to zero. These suggest emergence of nearly deconfined excitations. However, the splitting between the two Goldstone modes, caused by the altermagnetism, remains at the transition point, indicating that the transition is weakly first-order and proximate to a putative deconfined quantum critical point. Our results indicate that the altermagnetism of the system has an interesting interplay with the deconfined quantum criticality and provides viable means of detecting this exotic phase transition.
Acknowledgments.— We thank Runze Chi, Wenan Guo, Bruce Normand, Wei Li, Zhengxin Liu, Anders W. Sandvik, Yiming Wang, Jize Zhao for helpful discussions. This work is supported by the National Key R&D Program of China (Grant No. 2023YFA1406500), the National Natural Science Foundation of China (Grant Nos. 12334008, 12274458, 12174441).
SUPPLEMENTAL MATERIAL – Spin excitations of the Shastry-Sutherland model – altermagnetism and proximate deconfined quantum criticality
Hongyu Chen1, Guijing Duan1, Changle Liu2, Yi Cui1,3, Weiqiang Yu1,3, Z.Y.Xie1,3, Rong Yu1,3
1School of Physics and Beijing Key Laboratory of Opto-electronic Functional Materials and Micro-nano Devices, Renmin University of China, Beijing 100872, China
2School of Engineering, Dali University, Dali, Yunnan 671003, China
3Key Laboratory of Quantum State Construction and Manipulation (Ministry of Education), Renmin University of China, Beijing, 100872, China
I iPEPS method for calculating the ground state and excited states
Figure S-1: (a): The iPEPS setup in the calculation of the ground state of the SS model. The tensors are arranged at the sites of the SS lattice. The bright-colored sites indicate a $2\times 2$ unit cell used in the calculation. (b): Illustration of the single-mode iPEPS ansatz for the calculation of excited states. In each term of the summation, a single local tensor $A_{\alpha}$ located at site $r$ belonging to a specified sublattice is modified to a different tensor $B_{\alpha}$. Here $\alpha=1,2,3,4$, corresponding to a $2\times 2$ unit cell.
In this work, we apply the infinite projected entangled pair state (iPEPS) method [56, 57] to calculate the ground state of the Shastry-Sutherland (SS) model. A sketched iPEPS setup with a $2\times 2$ unit cell is shown in Fig. S-1(a). In the iPEPS setup, a local tensor is defined on each site of the SS lattice. As shown in Fig.  S-1, each local tensor has a physical index corresponding to the spin degrees of freedom, with dimension $2$ and denoted as curve, and four virtual indices corresponding to the four links of the site, with dimension $D$ referred to as the iPEPS bond dimension and denoted as straight. In this work, as emphasized as bright color in Fig. S-1, the unit cell is chosen to be $2\times 2$, and this is the minimal size compatible with the ground state of the SS model.
Given an iPEPS state $|\Psi\rangle$, the ground state energy can be evaluated according to the fundamental quantum mechanics
$E=\frac{\sum_{\langle ij\rangle}\langle\Psi|\hat{H}_{ij}(J)|\Psi\rangle}{\langle\Psi|\Psi\rangle}+\frac{\sum_{\langle\langle ij\rangle\rangle}\langle\Psi|\hat{H}_{ij}(J^{\prime})|\Psi\rangle}{\langle\Psi|\Psi\rangle}$ (S-1)
where $\langle ij\rangle$ and $\langle\langle ij\rangle\rangle$ denote nearest neighbors and second nearest neighbors, respectively. In this sense, the energy can be regarded as a complicated function of the local tensors, the variational parameters in the iPEPS ansatz. The denominators and numerators in Eq. (S-1) can be effectively calculated by using the corner transfer matrix renormalization group method [58, 59, 60, 75]. Then, starting from an arbitrary state, the ground state is obtained by variationally minimizing the energy through gradient-based optimization, e.g., the L-BFGS optimization strategy [76]. In this work, the energy gradients concerning the local tensors are obtained effectively by the automatic differentiation techniques [61, 62].
Once the ground state $|\Psi_{0}\rangle$ is obtained, we can construct the iPEPS ansatz for the excited state through the single-mode approximation [63, 64, 65]. More specifically, the excited ansatz with lattice momentum $k$ reads as
$|\Psi_{k}(B_{\alpha})\rangle=\sum_{r\in\alpha}e^{ikr}|\Psi_{r}(B_{\alpha})\rangle$ (S-2)
where $\alpha$ denotes the four sublattices corresponding to the $2\times 2$ unit cell, and the summation is over all the sites belonging to the $\alpha$-th sublattice. Here $|\Psi_{r}(B_{\alpha})\rangle$ is the perturbed $|\Psi_{0}\rangle$ state obtained by replacing the local tensor $A_{\alpha}$ at $r$ in $|\Psi_{0}\rangle$by a different tensor $B_{\alpha}$, which satisfies the orthogonal condition
$\langle\Psi_{r}(B_{\alpha})|\Psi_{0}\rangle=0.$ (S-3)
In practice, $B_{\alpha}$ can be chosen to be a vector in the null space of $E_{\alpha}$, which is the effective environment of $A_{\alpha}$ in $\langle\Psi_{0}|\Psi_{0}\rangle$ and satisfies $\langle\Psi_{0}|\Psi_{0}\rangle=\mathrm{Tr}E_{\alpha}A_{\alpha}$. For a given ground state with bond dimension $D$, we have $12D^{4}-$ 1 choices of $B_{\alpha}$ satisfying Eq.(S-3) for each $\alpha$, and thus $48D^{4}-$ 4 basis vectors, denoted as $|\Phi_{i}(k)\rangle$, in total to span the tangent space of the ground state. Then, we can calculate the matrix representation of the Hamiltonian in this non-orthogonal basis, and solving the generalized eigenvalue equation
$\langle\Psi_{i}(k)|\hat{H}|\Psi_{j}(k)\rangle=E(k)\langle\Psi_{i}(k)|\Psi_{j}(k)\rangle$ (S-4)
gives the excitation energies $E(k)$s and excited eigenstates $|\Phi_{i}(k)\rangle$s at momentum $k$ In this work, we have employed an effective derivative trick and a regularization procedure, introduced in Ref. [63, 64], to determine the matrices that appear in Eq. (S-4).
As long as the wave functions of the excited states $|\Phi_{i}(k)\rangle$s are obtained, we can use them as a complete set to calculate the spin dynamical structure factor (DSF) ($\alpha=x,y,z$)
[…]
$=$ $\sum_{n}\delta(E_{n}-E_{0}-\omega)w^{\alpha}_{n}(k)$
where $w^{\alpha}_{n}(k)\equiv|\langle\Phi_{n}(k)|S_{k}^{\alpha}|\Psi_{0}\rangle|^{2}$ is the spectral weight for each mode. If regarded as the overlap of two excited-state wave functions, the spectral weight can be tackled similarly by the derivative trick [64]. In practical calculations, a Lorentzian broadening is employed to deal with the delta function and mimic the finite temperature effect.
II Linear spin wave and bond-operator theories for the SS model
To better understand the TN results in the antiferromagnetic (AFM) and the plaquette phases, we study the spin excitations in these two phases by using the linear spin wave (LSW) and bond-operator theories, respectively.
II.1 Linear spin wave theory for spin excitations in the antiferromagnetic phase
Here we provide the LSW theory for spin excitations in the AFM phase of the SS model. We pick up a $2\times 2$ magnetic unit cell in the AFM state. Each spin on the SSL can be identified by the combination of a vector ${\mathbf{r}}$ labeling the unit cell position and a sublattice index $l$ ($l=1,2,3,4$) within each unit cell. We introduce a unit vector $\mathbf{n}_{l}$ to denote the orientation of the spin at sublattice $l$, and then define two additional unit vectors $\mathbf{u}_{l}$ and $\mathbf{v}_{l}$ according to ${\mathbf{u}_{l}\cdot\mathbf{n}_{l}=0}$ and ${\mathbf{v}_{l}=\mathbf{n}_{l}\times\mathbf{u}_{l}}$, namely, $\mathbf{n}_{l}$, $\mathbf{u}_{l}$ and $\mathbf{v}_{l}$ are orthogonal to each other. Next we perform the Holstein-Primakoff transformation for the spin operator $\mathbf{S}_{{\mathbf{r}}l}$,
$\mathbf{n}_{l}\cdot\mathbf{S}_{{\mathbf{r}}l}$ $=$ $S-b_{{\mathbf{r}}l}^{\dagger}b_{{\mathbf{r}}l},$ (S-6)
$(\mathbf{u}_{l}+i\mathbf{v}_{l})\cdot\mathbf{S}_{{\mathbf{r}}l}$ $=$ $({2S-b_{{\mathbf{r}}l}^{\dagger}b_{{\mathbf{r}}l}})^{\frac{1}{2}}b_{{\mathbf{r}}l},$ (S-7)
$(\mathbf{u}_{l}-i\mathbf{v}_{l})\cdot\mathbf{S}_{{\mathbf{r}}l}$ $=$ $b_{{\mathbf{r}}l}^{\dagger}({2S-b_{{\mathbf{r}}l}^{\dagger}b_{{\mathbf{r}}l}})^{\frac{1}{2}}.$ (S-8)
Due to the bipartite lattice structure, we adopt two sets of Fourier transforms. On the sublattice $l$ with an up spin:
$b_{{\mathbf{r}}l}=\sqrt{\frac{2}{N}}\sum_{\mathbf{k}\in{\text{MBZ}}}b_{\mathbf{k}l}e^{i\mathbf{R}_{{\mathbf{r}}l}\cdot\mathbf{k}}$ (S-9)
and on the sublattice $l^{\prime}$ with a down spin:
$b_{{\mathbf{r}}l^{\prime}}=\sqrt{\frac{2}{N}}\sum_{\mathbf{k}\in{\text{MBZ}}}b_{\mathbf{k}l^{\prime}}e^{-i\mathbf{R}_{{\mathbf{r}}l^{\prime}}\cdot\mathbf{k}}$ (S-10)
where $\mathbf{R}_{{\mathbf{r}}l}$ denotes the position of a spin on the SSL labeled by the magnetic unit cell position ${\mathbf{r}}$ and sublattice index $l$. The spin Hamiltonian can be rewritten in terms of boson bilinears as
$H_{\text{LSW}}=\sum_{\mathbf{k}\in{\text{MBZ}}}\Psi(\mathbf{k}){}^{\dagger}h(\mathbf{k})\Psi(\mathbf{k})+const.,$ (S-11)
where
$\Psi(\mathbf{k})=[b_{\mathbf{k}1},b_{\mathbf{k}2},b^{\dagger}_{\mathbf{k}3},b^{\dagger}_{\mathbf{\mathbf{k}}4}]^{T},$ (S-12)
and
$h(k)=\left(\begin{array}[]{cccc}2J-\frac{J^{\prime}}{2}&\frac{J^{\prime}}{2}e^{\frac{i(k_{x}-k_{y})}{2}}&J\cos(\frac{k_{x}}{2})&J\cos(\frac{k_{y}}{2})\\ \frac{J^{\prime}}{2}e^{\frac{-i(k_{x}-k_{y})}{2}}&2J-\frac{J^{\prime}}{2}&J\cos(\frac{k_{y}}{2})&J\cos(\frac{k_{x}}{2})\\ J\cos(\frac{k_{x}}{2})&J\cos(\frac{k_{y}}{2})&2J-\frac{J^{\prime}}{2}&\frac{J^{\prime}}{2}e^{\frac{-i(k_{x}+k_{y})}{2}}\\ J\cos(\frac{k_{y}}{2})&J\cos(\frac{k_{x}}{2})&\frac{J^{\prime}}{2}e^{\frac{i(k_{x}+k_{y})}{2}}&2J-\frac{J^{\prime}}{2}\end{array}\right).$ (S-13)
Then we can diagonalize $H_{\text{LSW}}$ via the Bogoliubov transformation with ${\Psi(\mathbf{k})=T_{\mathbf{k}}\Phi(\mathbf{k})}$, where
$\Phi(\mathbf{k})=[\beta_{\mathbf{k}1},\beta_{\mathbf{k}2},\beta_{\mathbf{k}3}^{\dagger},\beta_{\mathbf{k}4}^{\dagger}]^{T},$ (S-14)
is the diagonalized basis and $T_{\mathbf{k}}$ is the transformation matrix.
The general form of spin-wave dispersion is complex but can be easily computed numerically. Here we present the analytical form of the dispersions of the two Goldstone modes along the $\Gamma$-M direction near the $\Gamma$ (and equivalently, M) point. Up to the second-order terms in the wave vector $k_{x}=k_{y}$, the dispersions for $J\neq J^{\prime}$ are
$\begin{matrix}E_{+}&=\sqrt{\frac{2J^{2}(J-J^{\prime})}{2J-J^{\prime}}}k_{x}+\frac{J^{2}J^{\prime}}{2(2J-J^{\prime})^{2}}k_{x}^{2},\\ \\ E_{-}&=\sqrt{\frac{2J^{2}(J-J^{\prime})}{2J-J^{\prime}}}k_{x}-\frac{J^{2}J^{\prime}}{2(2J-J^{\prime})^{2}}k_{x}^{2},\end{matrix}$ (S-15)
respectively. Note that the coefficients of the linear-in-$k_{x}$ terms are the same, while the splitting caused by the altermagnetism appears in the subleading $k_{x}^{2}$ terms.
At $J=J^{\prime}$ where the AFM ground state becomes unstable in the LSW theory, these two magnon bands are softened with dispersion relations
$\begin{matrix}E_{+}&=\frac{Jk_{x}^{2}}{2},\\ \\ E_{-}&=0.\end{matrix}$ (S-16)
II.2 Bond operator theory for the triplet excitations in the plaquette phase
In the plaquette phase, the spin rotational symmetry preserves and the LSW theory is not applicable. To derive the low-lying triplet excitations in this phase, we project the Hamiltonian onto the low-energy subspace composed of the singlet ground state and the lowest triplet excitations.
Define a vacuum $|0\rangle$ state and four boson operators which yield the four physical states (one singlet and three triplets) by $|s\rangle=s^{\dagger}|0\rangle$, $|t_{\alpha}\rangle=t_{\alpha}^{\dagger}|0\rangle$, respectively [78, 77]. The projection operators are expressed $Z^{st_{\alpha}}=s^{\dagger}t_{\alpha}$, $Z^{t_{\alpha}t_{\beta}}=t_{\alpha}^{\dagger}t_{\beta}$. The spins represented via these boson operators are
$S_{l}^{\alpha}=\frac{(-1)^{l}}{\sqrt{6}}(s^{\dagger}t_{\alpha}+t_{\alpha}^{\dagger}s)-\frac{i}{4}\sum_{\beta,\gamma}\epsilon_{\alpha\beta\gamma}t_{\beta}^{\dagger}t_{\gamma}$ (S-17)
where $\alpha,\beta,\gamma=x,y,z$ denote the spin components and $l$ is the sublattice index $(l=1,2,3,4)$. Note that the Hilbert space of the bosons is larger than the original plaquette Hilbert space and includes unphysical states. To restrict the boson Hilbert space to its physical sector, a hard-core constraint must be imposed: $s^{\dagger}s+\sum_{\alpha}t_{\alpha}^{\dagger}t_{\alpha}=1$. By condensing the singlets $s^{\dagger}=s\approx\langle s\rangle$, we rewrite the Hamiltonian in terms of triplet operators
$H=H_{0}+\sum_{\mathbf{k}\in{\text{MBZ}}}\Psi(\mathbf{k}){}^{\dagger}h(\mathbf{k})\Psi(\mathbf{k})+const.,$ (S-18)
where $H_{0}=E_{s}s^{\dagger}s+\sum_{\alpha}E_{\alpha}t_{\alpha}^{\dagger}t_{\alpha}$ is the energy of a single plaquette and
$\Psi(\mathbf{k})=[t_{\mathbf{k}x},t_{\mathbf{k}y},t_{\mathbf{\mathbf{k}}z},t^{\dagger}_{\mathbf{-k}x},t^{\dagger}_{\mathbf{-k}y},t^{\dagger}_{\mathbf{\mathbf{-k}}z}]^{T}$ (S-19)
Then following the stantard Bogoliubov diagonalization we can derive the dispersion of the lowest triplet as
$E=\sqrt{J^{2}+\frac{2}{3}J(J^{\prime}-2J)(\cos(k_{x})+\cos(k_{y}))}.$ (S-20)
We compare the triplet dispersion with that calculated by using the TN method in Fig. 3 of the main text, and they agree well. The small difference is likely caused by the finite bond dimension $D$ in the TN calculation.
III Numerical results on the spin excitation spectra
Figure S-2: Total spin DSF ($\mathcal{S}^{xx+yy+zz}$) in the AFM phase of the SS model calculated by iPEPS method with $D=3$ (in (a) and (b)) and $D=2$ (in (c) and (d)), at $J/J^{\prime}=0.9$ (in (a) and (c)) and $J/J^{\prime}=1.1$ (in (b) and (d)).
Figure S-3: Finite $D$ scaling of the spin gap of the Goldstone modes (at M point) in the AFM phase of the SS model at $J/J^{\prime}=1.1$. The dashed line shows a best fit to the gap, which vanishes in the large $D$ limit.
Here we show numerical results on the spin excitation spectra in additional to those in the main text. The spectra illustrate the spin DSF In Fig. 2(b) we have shown the spectra in the AFM phase at $J/J^{\prime}=1.1$ calculated in the TN approach with bond dimension $D=3$. In Fig. S-2, we show calculated spectra for several other $J/J^{\prime}$ values and with different bond dimension $D$. One sees that the spectra show similar features with two split magnon bands, which are the Goldstone modes, at low energy and several longitudinal modes at high energy. Interestingly, the spin gap of the Higgs mode decreases with decreasing $J/J^{\prime}$ toward the plaquette-to-AFM transition, as discussed in the main text.
As an effect of finite $D$, the Goldstone modes exhibit a finite gap. We have calculated the spin gaps with other $D$ values. As shown in Fig. S-3, the spin gap reduces with increasing $D$ and eventually vanishes in the large $D$ limit. This indicates that the gapless feature of Goldstone modes is captured by the tensor network calculation in this limit.
Figure S-4: (a): Difference of the transverse DSF $\Delta\mathcal{S}=\mathcal{S}^{+-}(\mathbf{q},\omega)-\mathcal{S}^{-+}(\mathbf{q},\omega)$ in the AFM phase of the SS model at $J/J^{\prime}=1.1$ in the LSW theory. (b): Same as (a), but calculated by the tensor network method. In either case, the opposite sign of $\Delta\mathcal{S}$ reflects the opposite chirality of the magnon modes.
To show that the chiral magnon excitations in the AFM phase of the SS model, we calculate the difference of the transverse component of the DSF $\Delta\mathcal{S}=\mathcal{S}^{+-}(\mathbf{q},\omega)-\mathcal{S}^{-+}(\mathbf{q},\omega)$. Note that $S_{i}^{+}S_{j}^{-}-S_{i}^{-}S_{j}^{+}\propto(\mathbf{S}_{i}\times\mathbf{S}_{j})_{z}$, so that the sign of $\Delta\mathcal{S}$ is able to detect the chirality of the spin excitations. In Fig. S-4 we compare results of $\Delta\mathcal{S}$ calculated by LSW theory and tensor network method. Both results show a finite splitting between the two Goldstone modes with opposite sign of $\Delta\mathcal{S}$. Also note that the magnon bands along $\Gamma$-M and $\Gamma$-M also have opposite sign. Interestingly, we find that the optical magnon bands also exhibit opposite chirality. These results verify the existence of chiral magnons in the AFM state, which is caused by the symmetry of altermagnetism.
Figure S-5: (a): Density of states (DoS) with excitation energy at several $J/J^{\prime}$ values in the AFM phase. (b): Same as (a) but in the plaquette phase.
Besides the collective magnon excitations and the longitudinal modes, the tensor calculation also resolves continuum excitations. The spectral weights of these excitations are much smaller than the collective modes but as shown in Fig. S-5, they cause a broad peak in the density of states (DoS) at high energy. In the AFM phase, this broad peak appears in energies above the Higgs gap (above $\sim 2J^{\prime}$ at $J/J^{\prime}=1.1$), and the peak energy decreases with varying $J/J^{\prime}$ toward the plaquette-to-AFM transition. This behavior is consistent with the emergence of deconfined excitations although it is hard to tell whether the continuum has a spinon or magnon origin. In the plaquette phase, we can also identify continuous excitations at energies above the triplet gap (above $\sim J^{\prime}$ at $J/J^{\prime}=0.7$) in the DoS plots (Fig. S-5(b)). The behavior of energies of these excitations is similar to that in the AFM phase.
Figure S-6: (a): Field dependence of gaps of several low-energy excitations in the plaquette phase of the SS model, calculated by the TN method. The multiplets split into $S^{z}=+1,-1,0$ states. (b): Field dependence of low-energy excitations of a 4-spin cluster with nearest neighbor AFM interactions, calculated by the ED. The splitting indicates that the 7-fold multiplet contains two degenerate triplets and one singlet.
To understand the singlet and the 7-fold multiplet modes above the lowest triplet mode in the plaquette phase resolved by the tensor network calculation in the main text, we apply a small magnetic field to split the multiplets. We also compare our tensor network results with those from exact diagonalization on a four-spin plaquette with nearest neighbor Heisenberg interaction. As shown in Fig. S-6(a), the 7-fold quasi-degenerate excitations splits under a small field. The gaps of three modes are unchanged under the field, while the gaps of two modes increase with the field, gaps of another two modes decrease with the field. Note that the slops of the gaps are the same for these modes, and they are also same to the slops of the split lowest triplet excitations. These indicate that the 7-fold multiplet actually consists of one singlet and two degenerate triplets. This conclusion is further confirmed by our exact diagonalization calculation shown in Fig. S-6(b). At zero field, there are indeed 7-fold degenerate multiplet excitations above the triplet ones. With applying a small magnetic field, The 7-fold multiplet splits into a pair of triplets and one singlet. Note that the singlet excitation in between the lowest triplet and the 7-fold multiplet seen in the tensor network result in Fig. 3 of the main text does not show up in our 4-spin cluster exact diagonalization calculation. It is then interpreted as an inter-plaquette bound state.
Figure S-7: Evolution of gaps at the M point of the triplet mode (TM), Higgs mode (HM), and Goldstone modes (GM) with $J/J^{\prime}$ near the plaquette-to-AFM transition in the iPEPS calculation with $D=3$ (in (a)) and $D=5$ (in (b)). In either case we observe softening of gaps of both the triplet mode and Higgs mode when approaching the plaquette-to-AFM transition.
In Fig. 4(a) of the main text we discuss the evolution of triplet mode, Higgs mode, and Goldstone modes with varying $J/J^{\prime}$. It is found that the gaps of both triplet mode and Higgs mode drop rapidly to small values close to zero, and become quasi-degenerate with the Goldstone modes when approaching the plaquette-to-AFM transition. Here in Fig. S-7 we show results with several other bond dimension $D$ values. We observe the same behavior that the gaps of both triplet mode and Higgs mode are softened near the transition. This behavior strongly suggests emergence of nearly deconfined excitations at the plaquette-to-AFM transition.