### v2

# Spin excitations of the Shastry-Sutherland model – altermagnetism and deconfined quantum criticality

Abstract
Frustrated quantum magnets can host a variety of exotic spin excitations, including fractionalized spin excitations coupled to emergent gauge fields at deconfined quantum critical points (DQCPs) and chiral magnons in altermagnets. Here, we investigate the spin excitation spectra of the highly frustrated $S=1/2$ antiferromagnetic (AFM) Shastry-Sutherland model, focusing on the evolution of low-energy collective modes from the Néel AFM phase to the plaquette valence bond solid (PVBS). We demonstrate that the AFM state exhibits altermagnetic behavior, characterized by a non-relativistic splitting between two chiral magnon bands. Furthermore, we identify two additional low-energy excitations: a Higgs mode in the longitudinal excitation channel and an $S=0$ excitation with vanishing spectral weight. As the system approaches the AFM-to-PVBS transition, both these modes soften along with the lowest-energy triplet and singlet modes in the PVBS state. The closing gap of the Higgs mode, combined with the nearly degenerate velocities of $S=1$ and $S=0$ excitations, provides spectral evidence that the AFM-to-PVBS transition is proximate to a DQCP with emergent $O(4)$ symmetry. Our results help clarify the spectral signature of a broad class of symmetry enhanced quantum phase transitions including deconfined quantum criticality.
I Introduction
Symmetry provides an organizational principle to classify complex states of matter, and based on this, many exotic phenomena are understood TI_Review ; Senthil_AnnRevCondMattPhys_2015 ; Witten_NP_2018 . One prominent example is the recently proposed altermagnet, a collinear antiferromagnetic (AFM) state in which the two sublattices with oppositely polarized spins are connected by the rotational symmetry instead of inversion or translation Smejkal_PRX_2022A ; Smejkal_PRX_2022B . The distinct symmetry, which is organized by the spin space group theory Yang_arXiv:2105.12738 , causes a non-relativistic spin-dependent splitting of energy bands and may give rise to some exotic effects, such as anomalous Hall effect Smejkal_PRX_2022A ; Smejkal_NatRevMater7_2022 . Up to now, significant progress has been made in the study of itinerant systems Gonzalez_Betancourt_PRL_2023 ; Hariki_PRL_2024 ; Ferrari_arXiv:2408.00841 ; Wu_PRB_2007 ; Liu_NC_2021 ; Satoru_JPSJ_2019 . For insulating altermagnets, though a symmetry dictated splitting in magnon bands has been predicted Smejkal_PRL131_256703_2023 , existing studies are mostly limited to several model systems Ma_PRB110_064426_2024 ; Weissenhofer_PRB110_094427_2024 ; Yao_arXiv_2024 ; Zhaojize_arxiv .
Symmetry can also dictate phase transitions. In the Landau-Ginzburg-Wilson paradigm, a system transitions from a disordered state to an ordered one by spontaneously breaking a symmetry. With quantum fluctuations the transition can be continuous and occur at zero temperature, giving rise to intriguing quantum critical phenomena sachdev_2011 ; Coldea_2010 ; Senthil2004 ; Cui2023DQCP ; Cui2019SCVO ; Lee2019 ; E8_2021 ; Xu2022 ; agrawal2020quantum ; fey2019quantum ; zhang2011exploring . For a transition between two ordered phases, the Landau paradigm predicts a first-order transition. However, this scenario has recently been challenged. It was proposed that the transition between a valence bond solid (VBS) and an antiferromagnetic (AFM) phase can be continuous, e.g., via a deconfined quantum critical point (DQCP) Senthil2004 . At the DQCP, the establishment of one order is accompanied by the destruction of the other order via proliferation of topological excitations, giving rise to emergent deconfined fractionalized excitations as well as enhanced continuous symmetry among the seemingly unrelated two types of order parameters. This scenario has motivated extensive studies on DQCP Sandvik2007 ; Meng_EasyplaneDQCP ; Shao_Science ; Senthil_Vojta_Sachdev:2004 ; Liu_Vojta:2022 ; Guo_PRL_2024 . However, whether a DQCP can be realized in quantum spin models and what would be its spectral signatures are still challenging questions.
Tuning the frustration is a promising way to induce a VBS-to-AFM transition, and a prominent example is given by the Shastry-Sutherland (SS) model Shastry1981 . The model is defined on the SS lattice as illustrated in Fig. 1(a). It contains orthogonal spin dimers (with coupling $J^{\prime}$) connected by frustrated nearest neighbor (n.n.) bonds (with coupling $J$). With increasing $J/J^{\prime}$, the ground state experiences a series of quantum phase transitions from a dimer singlet (DS) phase to a plaquette valence bond solid (PVBS) at $J/J^{\prime}\approx 0.68$, then to a Néel AFM state for $J/J^{\prime}\sim 0.8$ (see Fig. 1(b)) Koga2000 ; Chung2001 ; Pixley2014 ; Corboz2013 ; Boos2019 ; Lee2019 ; yang2022quantum ; Sengupta_PRL:2013 ; Wang_PRL_2023 . Although the DS-to-PVBS transition is clearly strongly first-order, the nature of the PVBS-to-AFM transition remains elusive: Several tensor network calculations Corboz2013 ; Xi2023 showed it to be weakly first-order, while some other studies yang2022quantum ; Corboz_arXiv_2025 found an intervening gapless quantum spin liquid in between the PVBS and AFM phases. Interestingly, one DMRG work Lee2019 suggested the transition is via a DQCP with an emergent $O(4)$ symmetry, and quantum scaling behavior near this point has been revealed by a finite-size tensor network analysis WenyuanLiu_PRL_2024 . One prominent signature of a DQCP is the emergence of deconfined fractionalized spin excitations. Although spin excitations in both PVBS and AFM states have been extensively studied Mila_BOT ; Pinaki_PRB92_094440_2015 ; WangZ2018 ; Wang2023 , whether deconfined excitations emerge at the transition is still unclear.
The phase diagram of the SS model well describes the evolution of low-temperature phases of the quasi-2D antiferromagnet $\mathrm{SrCu}_{2}(\mathrm{BO}_{3})_{2}$ under pressure tuning Kageyama1999 ; PhysRevLett.82.3701 ; Koga2000 ; Waki2007 ; haravifard2016crystallization ; Zayed2017 ; Bettler2020 ; Guo2020 ; Jimenez2021 , giving the hope to realize a DQCP in this material. Remarkably, a recent NMR study in this compound Cui2023DQCP reported an unusually large anomalous dimension, $\eta\sim 0.2$, at a field-induced PVBS-to-AFM transition under pressure, implying behavior in proximity to a DQCP. However, direct spectral evidence on the deconfined excitations at zero field is absent.
In this article, we study spin excitations of the SS model by using the state-of-the-art tensor network technique for infinite lattice systems, i.e., the infinite projected entangled pair state (iPEPS) method. This prevents the severe finite-size effect in usual numerical calculations and offers a new perspective on the nature of emergent DQCP phenomena. Rather than directly searching for the elusive emergent deconfined excitation continuum at the DQCP, we investigate the evolution of the collective excitation modes (namely, Higgs and Goldstone modes, etc.) associated with the broken emergent enhanced symmetry near the transition in the ordered phases. We demonstrate that their simultaneous, significant softening directly points to the existence of a DQCP. In the AFM state, we find that the altermagnetic nature of the Néel order results in a finite splitting with opposite chirality in the magnon bands along the $\Gamma$ to M direction of the Brillouin zone. Besides the transverse Goldstone excitations, we have identified two additional low-energy modes: a Higgs mode in the longitudinal excitation channel and an $S=0$ mode with vanishing spectral weight. As the system approaches the PVBS-to-AFM transition, these modes in the AFM phase, as well as the lowest triplet and singlet modes in the plaquette phase, soften and exhibit gap closing behavior at the transition point. We further find the velocities of the singlet, magnon, and Higgs modes become degenerate at the transition point. This, together with the gap closing behavior aforementioned, strongly supports the presence of a DQCP with an emergent O(4) symmetry at the PVBS-to-AFM transition.
Figure 1: (a): Sketch of the SS lattice and the corresponding altermagnetism. The two sublattices of the Néel state are connected by a $C_{4}$ rotation of the lattice about the center of the empty plaquette (without the dimer bond). A time-reversal symmetry then recovers the Néel order. (b): Sketched phase diagram of the SS model with $J/J^{\prime}$. The $J/J^{\prime}$ values at the two transitions are taken from Ref. Xi2023 . (c): LSW dispersion in the AFM state of the SS model showing the chiral magnon bands with a non-zero splitting along the $\Gamma-M$ direction of the Brillouin zone. $\mathcal{L}$($\mathcal{R}$) refers to the left(right)-handed chiral magnon bands.
II Model and method
The Hamiltonian of the SS model reads as
$\mathcal{H}=J\sum_{\langle i,j\rangle}\mathbf{S}_{i}\cdot\mathbf{S}_{j}+J^{\prime}\sum_{\langle\langle i,j\rangle\rangle}\mathbf{S}_{i}\cdot\mathbf{S}_{j},$ (1)
where $\mathbf{S}_{i}$ is an $S=1/2$ spin operator defined on site $i$, $J$ and $J^{\prime}$ are AFM inter- and intra-dimer couplings on the SS lattice, and we set $J^{\prime}=1$ in this work. We apply the iPEPS method verstraete2004renormalization ; orus2014practical ; corboz2014competing ; orus2009simulation ; PhysRevB.84.041108 to calculate the ground states of the model. The iPEPS wavefunction is obtained by variationally minimizing the ground state energy using gradient-based optimization, where the gradients are calculated through the automatic differentiation technique SRGAD ; liao2019differentiable . Then we construct the iPEPS ansatz for excited states through the single-mode approximation, and solve the generalized eigenvalue problem for the Hamiltonian in the tangent space spanned by these excited states  ponsioen2020excitations ; ponsioen2022automatic ; chi2022spin . Details of the methods are explained in the Appendix. With the information of the excited states, we calculate the spin dynamical structure factor (DSF) $\mathcal{S}^{\alpha\beta}(\mathbf{q},\omega)=\frac{1}{N}\sum_{i,j}\int dte^{i\mathbf{q}\cdot(\mathbf{r}_{i}-\mathbf{r}_{j})}e^{i\omega t}\langle S_{i}^{\alpha}(t)S_{j}^{\beta}(0)\rangle$, where $\alpha,\beta$ refer to the spin components. The excitation spectra are calculated within the tensor network states up to bond dimension $D=5$. We also compare the results with linear spin wave (LSW) and bond-operator theories, which are introduced in detail in the Appendix.
III Altermagnetism and chiral magnons in the SS model
The space group of the SS lattice is non-symmophic and the two sublattices of the Néel AFM state are connected by neither translational nor inversion, but a $C_{4}$ rotation about the center of the empty plaquette. The AFM order then recovers by further applying a time-reversal $\mathcal{T}$ symmetry to the spins (see Fig. 1(a)). This indicates that the Néel AFM state on the SS lattice is an altermagnet with a $d$-wave symmetry. The spin space group symmetry dictates a non-zero splitting with opposite chirality of the magnon bands along the diagonal directions of the Brillouin zone. This is confirmed by the LSW calculation shown in Fig. 1(c). Along the $\Gamma-M$ direction, the acoustic (Goldstone) modes split into two branches with left- and right-handed chiralities, respectively. The splitting between the two chiral modes along $\Gamma$-M and $\Gamma$-M directions has opposite sign, and its magnitude reaches a maximum at the mid-point $\Sigma$ (and $\Sigma^{\prime}$). On the other hand, the two modes are degenerate along the $\Gamma$-X (and X-M) directions (Fig. 2(a)). These properties verify that the altermagnet has a $d$-wave nature. Note that the optical modes also split, but one band is completely flat along the $\Gamma-M$ due to frustration and has zero spectral weight.
Figure 2: (a): Spin DSF in the AFM phase of the SS model calculated by the LSW theory. The inset shows high symmetric directions of the Brillouin zone. (b): Total spin DSF ($\mathcal{S}^{xx+yy+zz}$) in the AFM phase calculated by the iPEPS method with $D=3$ at $J/J^{\prime}=1.1$. (c): The transverse component ($\mathcal{S}^{xx+yy}$) of the results in (b). (d): The longitudinal component ($\mathcal{S}^{zz}$) of the results in (b). The red, black, and white lines with dots in (b) and (d) label the magnon, Higgs, and $S=0$ modes, respectively.
IV Spin excitation spectra
We implement tensor network calculation to go beyond the LSW results on the spin excitation spectra. The calculated spectrum at $J/J^{\prime}=1.1$ is shown in Fig. 2(b). It exhibits richer structures than the LSW results in Fig. 2(a). To understand these structures, we calculate the transverse and longitudinal components of the DSF, $\mathcal{S}^{xx+yy}(\mathbf{q},\omega)$ and $\mathcal{S}^{zz}(\mathbf{q},\omega)$, and show them in Fig. 2(c) and (d), respectively. The transverse DSF resembles the LSW spectrum with three dispersive magnon bands. The finite gap at M (and $\Gamma$) of the Goldstone modes is an effect of finite bond dimension $D$ of the iPEPS ansatz, and vanishes in the large-$D$ limit (Fig.7 of the Appendix). One prominent feature is the splitting of Goldstone modes along the $\Gamma$-M direction. The opposite sign of $\Delta\mathcal{S}=\mathcal{S}^{+-}(\mathbf{q},\omega)-\mathcal{S}^{-+}(\mathbf{q},\omega)$ for the two modes shown in Fig. 8 justifies their opposite chirality, as expected for an altermagnet.
The spectrum of longitudinal modes, shown in Fig. 2(d), also consists of multiple bands, owing to the complex SS lattice structure (see discussion below). We can identify a mode (labeled by the black dashed line) that develops a strong resonance at the M point. The excitation gap of this mode decreases with reducing $J/J^{\prime}$ (see Fig.6 of the Appendix). We then attribute it to the Higgs mode which reflects the amplitude fluctuations of the AFM order parameter. In a two-dimensional AFM state, the Higgs mode is usually hardly visible because the gapless Goldstone modes give rise to a divergent longitudinal susceptibility and the Higgs mode can decay into a pair of Goldstone modes Poldolsky_PRB84_174522_2011 . Here it shows up in the tensor network calculation for the following reasons: The gapped Goldstone modes at finite $D$ remove the divergence, and the single-mode approximation adopted suppresses the two-magnon continuum but preserves the two-magnon bound state, which is just the Higgs mode. Just slightly above the Higgs mode, we find another excitation mode that is quasi-degenerate to the Higgs one at M point but with vanishing spectral weight. The gap of this mode follows a trend similar to that of the Higgs mode as $J/J^{\prime}$ decreases. We therefore identify this mode as a singlet ($S=0$) excitation originating from fluctuations of the plaquette VBS order parameter. Besides the collective excitation modes discussed above, the spectrum also contains a continuum, as shown in the broad peak above about $2J^{\prime}$ in the density of states in Fig. 9 of the Appendix. Interestingly, the peak moves to lower energy when reducing $J/J^{\prime}$ toward the PVBS-to-AFM transition, in a way similar to the behavior of the Higgs mode.
Figure 3: (a): Spin DSF in the plaquette phase calculated by the iPEPS method with $D=3$ at $J/J^{\prime}=0.7$. The white line with dots shows the lowest singlet excitation. The magenta dashed line labels the dispersion of the triplet excitations in the bond-operator theory. (b): $J/J^{\prime}$ dependence of extracted gaps at the M point of the Brillouin zone for several low-energy excitation modes in the plaquette phase, including, from bottom to top, a triplet (in red), a singlet (in black), and a 7-fold multiplet (in blue). The results are obtained by the iPEPS method with $D=5$. The dashed line shows the gap of the triplet excitations in the bond-operator theory.
For a comprehensive understanding, we also study the excitations in the plaquette phase, and a typical spectrum at $J/J^{\prime}=0.7$ is shown in Fig. 3(a). The lowest-energy excitations are found to be a triplet mode, whose dispersion can be described by the bond-operator theory Mila_BOT ; SM . Above the triplet mode, we can identify a singlet excitation which is almost dispersionless and has vanishing spectral weight (white dotted line in Fig. 3(a)). Fig. 3(b) illustrates excitation gaps at the M point of all identified low-energy modes, which, from bottom to top, include the triplet mode, the singlet mode, and a 7-fold quasi-degenerate multiplet. Note that the minor splitting within each triplet comes from numerical accuracy and can be eliminated by enforcing the spin SU(2) symmetry in the calculation. The gap of the triplet mode decreases with increasing $J/J^{\prime}$ toward the PVBS-to-AFM transition and the gap size is comparable with that from the bond-operator theory. Similar to the triplet, the lowest singlet excitation, which is interpreted as PVBS fluctuation, is also softened when approaching the transition. But the 7-fold multiplet does not exhibit significant softening. We calculate the quantum numbers of the 7-fold multiplet and find that this multiplet consists of two triplets and one singlet, which are the eigenstates of a single plaquette(see Fig. 10 of the Appendix). The inter-plaquette interaction then couples multiplets in each plaquette into bands that can be understood by a generalized bond-operator theory Pinaki_PRB92_094440_2015 . Note that each triplet excitation will further split to transverse and longitudinal modes once the magnetic order breaks the spin SU(2) symmetry, accounting for the complex multiband structure shown in Fig. 2 in the AFM phase. Above these low-energy excitations, the spectrum becomes continuous above about $J^{\prime}$ (Fig. 9 of the Appendix), similar to the case of the AFM phase.
Figure 4: (a): Evolution of low-energy collective (Higgs and Goldstone) modes across a DQCP with emergent O(4) symmetry. In the PVBS phase, the singlet (S) and triplet (T) modes serve as the Higgs and Goldstone modes associated with the emergent O(4) symmetry. In the AFM phase the Higgs mode (H) turns to a spin longitudinal mode (in black) whereas the singlet (S) and two magnon (M) modes constitute as the Goldstone modes. The DQCP is characterized by simultaneous gap closing of all four modes. (b): Similar to (a), but across a first-order transition with emergent O(4) symmetry. Contrast to the DQCP scenario, the gap of the Higgs mode keeps to be finite across the first-order transition. (c): Gaps at the M point of the triplet (T) and singlet (S) modes in the plaquette phase and of the magnons (M), the Higgs mode (H), and the singlet mode (S) with $J/J^{\prime}$, obtained in the iPEPS calculation with $D=5$. A simultaneous softening of all modes with gaps close to the magnon gap takes place near the PVBS-to-AFM transition at $(J/J^{\prime})_{c}\approx 0.79$. (d): Finite-$D$ analysis on the crossing points of the triplet, Higgs, and magnon modes. In the large-$D$ limit, all crossing points are close and can be extrapolated to $(J/J^{\prime})_{c}\approx 0.79$, suggesting the existence of a DQCP. (e): Finite $D$ analysis of the Higgs gap, $\Delta$, at the crossing point with the magnon mode, showing the closure of the Higgs gap in the large $D$ limit. (f): Evolution of the velocities $c$ with $J/J^{\prime}$ for the singlet mode in the PVBS phase together with the Higgs and magnon modes in the AFM phases. $c$ of the these modes coincide at $(J/J^{\prime})_{c}$. For each mode, $c$ is extracted by fitting the dispersion along the M to X direction with $\omega(\mathbf{q})=\sqrt{\Delta^{2}+c^{2}q^{2}}$.
V Probing the DQCP
With excitations studied in both PVBS and AFM phases, we now investigate the PVBS-to-AFM transition. Though a key characteristic of the DQCP is the emergence of deconfined fractionalized spin excitations, detecting them numerically is, in practice, challenging. This is because they only show up at the critical point and manifest as an elusive broad continuum in the spectrum. Here we adopt a different strategy to detect the possible DQCP by investigating the evolution of low-energy collective modes, i.e. Higgs and Goldstone modes associated with the broken emergent O(4) symmetry, in the PVBS and AFM phases.
Let us denote the three components of the AFM order parameter as $n_{1}$, $n_{2}$, and $n_{3}$, and the PVBS order parameter as $n_{4}$. In the SS model, they are defined as
$n_{\alpha=1,2,3}=\sum_{i}(-1)^{i}\langle S_{i}^{\alpha}\rangle,$ (2)
$n_{4}=\sum_{i}\left[(-1)^{i_{x}}\langle\mathbf{S}_{i}\mathbf{S}_{i+\hat{x}}\rangle+(-1)^{i_{y}}\langle\mathbf{S}_{i}\mathbf{S}_{i+\hat{y}}\rangle\right$]. (3)
If a DQCP exists, these four components should be combined to a 4-dimensional supervector by the emergent O(4) symmetry at the DQCP. When the system is driven to either the PVBS or AFM phase (with either $n_{4}\neq 0$ or $n_{3}\neq 0$) under an infinitesimal perturbation, the emergent O(4) symmetry is broken to O(3). This results in three Goldstone modes and one Higgs mode in each phase. In the AFM phase the Higgs mode is in the longitudinal spin ($S^{z}$) channel, whereas in the PVBS phase it corresponds to the amplitude fluctuations of $n_{4}$ which is a spin singlet. The Goldstone modes are created by the generators of the O(4) algebra that rotate the ordered component to other three disordered components. In the PVBS phase, $n_{4}$ is ordered, and $n_{1}$, $n_{2}$, and $n_{3}$ follow the fundamental spin SO(3) symmetry. As a result, the three Goldstone modes are just the lowest spin triplet excitations. Note that the triplet excitations are generally gapped since the PVBS order breaks discrete lattice symmetry, and strictly speaking, they are pseudo-Goldstone modes (in the context of emergent O(4) symmetry). But as the system approaches the critical point, they should exhibit a gap closing behavior and develop as the emergent gapless Goldstone modes. In the AFM phase, $n_{3}$ is ordered and in the vicinity of the critical point, $n_{1}$, $n_{2}$, and $n_{4}$ should follow an emergent O(3) symmetry. Besides the two magnons, which are Goldstone modes associated with $n_{1}$ and $n_{2}$, respectively, there is an additional emergent Goldstone mode associated with $n_{4}$, which is a spin singlet with a gap closure at the DQCP.
The evolution of these collective modes across an O(4) DQCP is sketched in Fig. 4(a). Approaching the critical point, three Goldstone modes and a Higgs mode soften, and their energies all go to zero at the DQCP. Given the O(4) symmetry at the DQCP, their difference disappears and they turn to 4-fold degenerate critical modes forming the lower edge of the continuum of deconfined spinons. This scenario contrasts to that of a first-order transition with an emergent O(4) symmetry. In the latter case, the composite 4-component order parameter is non-zero at the transition point. This breaks the O(4) symmetry in the ground state, giving rise to three emergent Goldstone modes and a gapped Higgs mode as illustrated in Fig. 4(b).
To examine which of the above two scenarios applies to the SS model, we have extracted gaps of the low-energy collective excitations at the M point, and their evolution across the PVBS-to-AFM transition for $D=5$ is plotted in Fig. 4(c). The gaps for other $D$ values are presented in Fig. 11 of the Appendix. The magnon gap is almost independent of $J/J^{\prime}$, confirming that it is caused by finite $D$, which closes in the entire AFM phase in the large $D$ limit (see Fig. 7 of the Appendix). On the other hand, the gaps of all other excitations, including the Higgs and singlet modes in the AFM phase and the triplet and singlet modes in the plaquette phase, drop rapidly and are close to the magnon gap when approaching the transition point $(J/J^{\prime})_{c}\approx 0.79\pm 0.02$. As shown in Fig. 4(d), the crossing points of these modes are all close and can be extrapolated to $(J/J^{\prime})_{c}$ in the large-$D$ limit. This implies that there is a single transition at $(J/J^{\prime})_{c}$ with gap closure of all excitations. We then plot the gap of the Higgs mode at the crossing point with magnons in Fig. 4(e). The Higgs gap indeed closes in the large $D$ limit, corroborating the DQCP scenario.
In Fig. 4(f) we further plot the extracted velocities $c$ of several excitation modes by fitting their dispersions along the M to X direction with $\omega(\mathbf{q})=\sqrt{\Delta^{2}+c^{2}q^{2}}$. Besides the gap closure behavior, we find the velocities of the singlet mode in the plaquette phase and the magnon and Higgs modes in the AFM phase coincide at $(J/J^{\prime})_{c}$. This once again supports the transition is via a DQCP: At the critical point, both Higgs and Goldstone modes become degenerate and appear as the lower edge of the deconfined spinon continuum $\omega_{\text{min}}(\mathbf{q})\sim c|\mathbf{q}|$, where the velocity $c$ is universal for all critical modes due to the emergent Lorentz invariance. The degeneracy of 4 modes – 1 from the $S=0$ (singlet) and 3 from the $S=1$ (Higgs and magnons) channel – supports the emergent symmetry of the DQCP is O(4).
VI Discussions and conclusions
The appearance of 4 degenerate low-lying excitation modes in both $S=0$ and $S=1$ channels when approaching the transition from either side as well as the gap closure of the Higgs mode provide direct evidence that the transition goes beyond the conventional O(3) Wilson-Fisher universality or a symmetry-enhanced first-order transition, but is instead via a DQCP with emergent O(4) symmetry. Quantum scaling behavior with novel critical exponents characterizing this DQCP is then expected WenyuanLiu_PRL_2024 . Some recent studies yang2022quantum ; Corboz_arXiv_2025 suggest that a quantum spin liquid can be stabilized in between the PVBS and AFM phases. Our analysis on the spin excitations cannot rule out a spin liquid phase, yet implies a spin liquid could only stabilized within a narrow parameter regime $0.77\lesssim J/J^{\prime}\lesssim 0.81$. Our calculation suggests the 4 critical modes keeps to be the only low-lying excitations within this parameter regime. Such a few number of low-lying excitations rules out the possibility of U(1) spin liquids Song_NC_2019 , and the putative spin liquid is most likely a $Z_{2}$ Dirac spin liquid Maity_arXiv_2025 , in which the $S=0$ excitation turns out to be the vison and opens a small gap while the other three modes represent gapless excitations in the spinon sector.
Moreover, our scenario naturally applies to a large class of models including the SS model, the checkerboard $J$-$Q$ model, and their variants. The PVBS-to-AFM transition in the checkerboard $J$-$Q$ model is found to be first-order with an enhanced O(4) symmetry. The low-lying excitations can then be described by the proximate DQCP scenario shown in Fig. 4(b). While the Higgs mode is gapped at the transition, the emergent Goldstone modes cause a gap closure and are responsible for the observed quantum scaling behavior Zhao_NP_2018 ; ChengchenLi_JPCM_2024 . Recent studies proposed a global phase diagram where the DQCP plays as a fine tuned multicritical point connecting the PVBS, AFM, and quantum spin liquid phases Zhao_PRL_2020 ; Lu_PRB_2021 . Our calculation suggests that the PVBS-to-AFM transitions in the SS model is either at or very close to the multicritical point whereas the one in the checkerboard $J$-$Q$ model is at a different regime of this proposed phase diagram, via a symmetry enhanced weakly first-order transition. Note that some other studies suggest the underlying conformal field theory (CFT) for the SO(5) DQCP could be non-unitary ma2020theory ; nahum2020note ; zhou2024so . Given that the O(4) DQCP is obtained by deforming the SO(5) CFT Lee2019 , this implies that realization of an O(4) DQCP on a lattice model without fine tuning could be delicate, which is reconciled with the proposed multicriticality picture.
In either the SS or the checkerboard $J$-$Q$ model, $C_{4}\mathcal{T}$ symmetry dictates the Néel AFM state of the system to be a $d$-wave altermagnet with split magnon bands. One might worry that the splitting of magnon bands could break the emergent Lorentz invariance and thereby turn the PVBS-to-AFM transition into a weakly first-order one. However, the magnon dispersion reads $\omega(\mathbf{q})=cq\pm wq_{x}q_{y}$, where the splitting is caused by the anisotropic $w$ term in the $q^{2}$ order. The magnon bands are still degenerate up to the leading linear-in-$q$ term. Therefore, the splitting, though a symmetry-breaking perturbation, is irrelevant in the low-energy effective theory of DQCP. Nonetheless, it provides a useful probe for the PVBS-to-AFM transition by signaling the symmetry breaking characteristic of an altermagnet.
We then discuss the implications of our results for experiments on the SS material SrCu2(BO3)2. A recent specific heat measurement find a weakly first-order PVBS-to-AFM transition at zero field LilingSun_arXiv_2024 , and the quantum scaling of the NMR spin-lattice relaxation rate in a magnetic field suggests the transition is proximate to a DQCP Cui2023DQCP . It is likely that the subleading interactions in the material, which goes beyond the description of the SS model, such as the Dzyaloshinsky-Moriya interaction or the interlayer coupling, drive the transition to first-order but still proximate to a putative DQCP. This well explains the unusually large anomalous dimension $\eta$ observed in the NMR measurement Cui2023DQCP .
Besides the possible DQCP, another unsettled issue is the nature of the PVBS state in experiment. Recent inelastic neutron scattering (INS) Zayed2017 and NMR Cui2023DQCP measurements suggest a full plaquette phase where the spin singlets are located on plaquettes with $J^{\prime}$ dimers. But most numerical studies (including ours) on the SS model found an empty plaquette (singlets on plaquettes without the $J^{\prime}$ dimers) ground state. Interestingly, for the plaquette state at $J/J^{\prime}=0.7$, if taking $J^{\prime}\sim 4$ meV estimated from the experiment, the calculated spin gaps of the lowest triplet mode and the 7-fold degenerate mode (including two degenerate triplets) are about 1 meV and 2 meV, which are consistent with the gap values of the lowest two excitations observed in the INS experiment Zayed2017 . It would be important to further compare the theoretical and experimental results on the dispersions and evolution of the spin gaps with pressure.
We have examined the lattice structure of the SS model material SrCu2(BO3)2, and find that it indeed supports an altermagnet even when the non-magnetic ions are included. To experimentally probe the altermagnetism, one may look for the splitting of the two magnon bands (at the order of $J^{\prime}$ near the $\Sigma$ point) in the spectrum of INS. The chiral magnons can be detected by resonant inelastic X-ray scattering (RIXS). Moreover, the transverse spin current induced by the chiral magnons will cause sizable spin Nernst effect in the AFM state Smejkal_PRX_2022B ; Cui_PRB_2023 when applying a temperature gradient along the non-splitting direction. Given that the Dzyaloshinsky-Moriya interaction is much smaller than $J$ and $J^{\prime}$ in SrCu2(BO3)2, the leading contribution should be attributed to the altermagnetism. Given that the splitting of the magnon bands in the altermagnetic phase is directly related to the O(4) symmetry breaking in the magnetic sector, the proposed experiments above, provide viable ways to detect the proximate deconfined quantum criticality.
In conclusion, we study the spin excitations of the SS model by using advanced tensor network method and examine the evolution of low-energy collective modes across the transition between the plaquette valence bond solid and the Néel antiferromagnetic phase. We show that the antiferromagnetic state of the model is an altermagnet supporting two split magnons (Goldstone modes) with opposite chirality. In addition, we have also identified a Higgs mode in the longitudinal spin channel and a spin singlet mode with vanishing spectral weight. The low-energy excitations in the plaquette valence bond solid contain a triplet mode and a singlet one. As the system approaches the transition, all low-energy collective excitations soften and their gaps close. Simultaneously, their velocities coalesce. These observations demonstrate that the transition is governed by a universal deconfined quantum criticality with an emergent O(4) symmetry. Our results clarify the spectral signatures in the vicinity of a deconfined quantum critical point and provide viable means of detecting this exotic phase transition.
Acknowledgements.
We thank Runze Chi, Wenan Guo, Bruce Normand, Wei Li, Zhengxin Liu, Anders W. Sandvik, Yiming Wang, Jize Zhao for helpful discussions. This work is supported by the National Key R&D Program of China (Grant No. 2023YFA1406500), the National Natural Science Foundation of China (Grant Nos. 12334008, 12274458, 12174441).
Appendix A iPEPS method for calculating the ground state and excited states
Figure 5: (a): The iPEPS setup in the calculation of the ground state of the SS model. The tensors are arranged at the sites of the SS lattice. The bright-colored sites indicate a $2\times 2$ unit cell used in the calculation. (b): Illustration of the single-mode iPEPS ansatz for the calculation of excited states. In each term of the summation, a single local tensor $A_{\alpha}$ located at site $r$ belonging to a specified sublattice is modified to a different tensor $B_{\alpha}$. Here $\alpha=1,2,3,4$, corresponding to a $2\times 2$ unit cell.
In this work, we apply the infinite projected entangled pair state (iPEPS) method verstraete2004renormalization ; orus2014practical to calculate the ground state of the Shastry-Sutherland (SS) model. A sketched iPEPS setup with a $2\times 2$ unit cell is shown in Fig. 5(a). In the iPEPS setup, a local tensor is defined on each site of the SS lattice. As shown in Fig.  5, each local tensor has a physical index corresponding to the spin degrees of freedom, with dimension $2$ and denoted as curve, and four virtual indices corresponding to the four links of the site, with dimension $D$ referred to as the iPEPS bond dimension and denoted as straight. In this work, as emphasized as bright color in Fig. 5, the unit cell is chosen to be $2\times 2$, and this is the minimal size compatible with the ground state of the SS model.
Given an iPEPS state $|\Psi\rangle$, the ground state energy can be evaluated according to the fundamental quantum mechanics
$E$ $=$ $\frac{\sum_{\langle ij\rangle}\langle\Psi|\hat{H}_{ij}(J)|\Psi\rangle}{\langle\Psi|\Psi\rangle}$ (4)
$+$ $\frac{\sum_{\langle\langle ij\rangle\rangle}\langle\Psi|\hat{H}_{ij}(J^{\prime})|\Psi\rangle}{\langle\Psi|\Psi\rangle}$
where $\langle ij\rangle$ and $\langle\langle ij\rangle\rangle$ denote nearest neighbors and second nearest neighbors, respectively. In this sense, the energy can be regarded as a complicated function of the local tensors, the variational parameters in the iPEPS ansatz. The denominators and numerators in Eq. (4) can be effectively calculated by using the corner transfer matrix renormalization group method corboz2014competing ; orus2009simulation ; PhysRevB.84.041108 ; VCTMRG2022 . Then, starting from an arbitrary state, the ground state is obtained by variationally minimizing the energy through gradient-based optimization, e.g., the L-BFGS optimization strategy LBFGS . In this work, the energy gradients concerning the local tensors are obtained effectively by the automatic differentiation techniques SRGAD ; liao2019differentiable .
Once the ground state $|\Psi_{0}\rangle$ is obtained, we can construct the iPEPS ansatz for the excited state through the single-mode approximation ponsioen2020excitations ; ponsioen2022automatic ; chi2022spin . More specifically, the excited ansatz with lattice momentum $k$ reads as
$|\Psi_{k}(B_{\alpha})\rangle=\sum_{r\in\alpha}e^{ikr}|\Psi_{r}(B_{\alpha})\rangle$ (5)
where $\alpha$ denotes the four sublattices corresponding to the $2\times 2$ unit cell, and the summation is over all the sites belonging to the $\alpha$-th sublattice. Here $|\Psi_{r}(B_{\alpha})\rangle$ is the perturbed $|\Psi_{0}\rangle$ state obtained by replacing the local tensor $A_{\alpha}$ at $r$ in $|\Psi_{0}\rangleby$ a different tensor $B_{\alpha}$, which satisfies the orthogonal condition
$\langle\Psi_{r}(B_{\alpha})|\Psi_{0}\rangle=0.$ (6)
In practice, $B_{\alpha}$ can be chosen to be a vector in the null space of $E_{\alpha}$, which is the effective environment of $A_{\alpha}$ in $\langle\Psi_{0}|\Psi_{0}\rangle$ and satisfies $\langle\Psi_{0}|\Psi_{0}\rangle=\mathrm{Tr}E_{\alpha}A_{\alpha}$. For a given ground state with bond dimension $D$, we have $2D^{4}-1$ choices of $B_{\alpha}$ satisfying Eq.(6) for each $\alpha$, and thus $8D^{4}-4$ basis vectors, denoted as $|\Phi_{i}(k)\rangle$, in total to span the tangent space of the ground state. Then, we can calculate the matrix representation of the Hamiltonian in this non-orthogonal basis, and solving the generalized eigenvalue equation
$\langle\Psi_{i}(k)|\hat{H}|\Psi_{j}(k)\rangle=E(k)\langle\Psi_{i}(k)|\Psi_{j}(k)\rangle$ (7)
gives the excitation energies $E(k)$s and excited eigenstates $|\Phi_{i}(k)\rangles$ at momentum $k$ . In this work, we have employed an effective derivative trick and a regularization procedure, introduced in Ref. ponsioen2020excitations ; ponsioen2022automatic , to determine the matrices that appear in Eq. (7).
As long as the wave functions of the excited states $|\Phi_{i}(k)\rangles$ are obtained, we can use them as a complete set to calculate the spin dynamical structure factor (DSF) ($\alpha=x,y,z$)
$\mathcal{S}^{\alpha\alpha}(k,\omega)$ $=$ $\int dte^{-i\omega t}\langle\Psi_{0}|S_{-k}^{\alpha}S^{\alpha}_{k}(t)|\Psi_{0}\rangle$ (8)
$=$ $\sum_{mn}\int dte^{-i(\omega+E_{0})t}\langle\Psi_{0}|S_{-k}^{\alpha}|\Phi_{m}(k)\rangle$
$\cdot$ $\langle\Phi_{m}(k)|e^{iHt}|\Phi_{n}(k)\rangle\langle\Phi_{n}(k)|S^{\alpha}_{k}|\Psi_{0}\rangle$
$=$ $\sum_{n}\delta(E_{n}-E_{0}-\omega)w^{\alpha}_{n}(k)$
where $w^{\alpha}_{n}(k)\equiv|\langle\Phi_{n}(k)|S_{k}^{\alpha}|\Psi_{0}\rangle|^{2}$ is the spectral weight for each mode. If regarded as the overlap of two excited-state wave functions, the spectral weight can be tackled similarly by the derivative trick ponsioen2022automatic . In practical calculations, a Lorentzian broadening is employed to deal with the delta function and mimic the finite temperature effect.
Appendix B Linear spin wave and bond-operator theories for the SS model
To better understand the tensor network results in the antiferromagnetic (AFM) and the plaquette phases, we study the spin excitations in these two phases by using the linear spin wave (LSW) and bond-operator theories, respectively.
B.1 Linear spin wave theory for spin excitations in the antiferromagnetic phase
Here we provide the LSW theory for spin excitations in the AFM phase of the SS model. We pick up a $2\times 2$ magnetic unit cell in the AFM state. Each spin on the SSL can be identified by the combination of a vector $\mathbf{r}$} labeling the unit cell position and a sublattice index $l$ ($l=1,2,3,4$) within each unit cell. We introduce a unit vector $\mathbf{n}_{l}$ to denote the orientation of the spin at sublattice $l$, and then define two additional unit vectors $\mathbf{u}_{l}$ and $\mathbf{v}_{l}$ according to $\mathbf{u}_{l}\cdot\mathbf{n}_{l}=0$} and $\mathbf{v}_{l}=\mathbf{n}_{l}\times\mathbf{u}_{l}$}, namely, $\mathbf{n}_{l}$, $\mathbf{u}_{l}$ and $\mathbf{v}_{l}$ are orthogonal to each other. Next we perform the Holstein-Primakoff transformation for the spin operator $\mathbf{S}_{{\mathbf{r}}l}$,
$\mathbf{n}_{l}\cdot\mathbf{S}_{{\mathbf{r}}l}$ $=$ $S-b_{{\mathbf{r}}l}^{\dagger}b_{{\mathbf{r}}l},$ (9)
$(\mathbf{u}_{l}+i\mathbf{v}_{l})\cdot\mathbf{S}_{{\mathbf{r}}l}$ $=$ $({2S-b_{{\mathbf{r}}l}^{\dagger}b_{{\mathbf{r}}l}})^{\frac{1}{2}}b_{{\mathbf{r}}l},$ (10)
$(\mathbf{u}_{l}-i\mathbf{v}_{l})\cdot\mathbf{S}_{{\mathbf{r}}l}$ $=$ $b_{{\mathbf{r}}l}^{\dagger}({2S-b_{{\mathbf{r}}l}^{\dagger}b_{{\mathbf{r}}l}})^{\frac{1}{2}}.$ (11)
Due to the bipartite lattice structure, we adopt two sets of Fourier transforms. On the sublattice $l$ with an up spin:
$b_{{\mathbf{r}}l}=\sqrt{\frac{2}{N}}\sum_{\mathbf{k}\in{\text{MBZ}}}b_{\mathbf{k}l}e^{i\mathbf{R}_{{\mathbf{r}}l}\cdot\mathbf{k}}$ (12)
and on the sublattice $l^{\prime}$ with a down spin:
$b_{{\mathbf{r}}l^{\prime}}=\sqrt{\frac{2}{N}}\sum_{\mathbf{k}\in{\text{MBZ}}}b_{\mathbf{k}l^{\prime}}e^{-i\mathbf{R}_{{\mathbf{r}}l^{\prime}}\cdot\mathbf{k}}$ (13)
where $\mathbf{R}_{{\mathbf{r}}l}$ denotes the position of a spin on the SSL labeled by the magnetic unit cell position $\mathbf{r}$} and sublattice index $l$. The spin Hamiltonian can be rewritten in terms of boson bilinears as
$H_{\text{LSW}}=\sum_{\mathbf{k}\in{\text{MBZ}}}\Psi(\mathbf{k}){}^{\dagger}h(\mathbf{k})\Psi(\mathbf{k})+const.,$ (14)
where
$\Psi(\mathbf{k})=[b_{\mathbf{k}1},b_{\mathbf{k}2},b^{\dagger}_{\mathbf{k}3},b^{\dagger}_{\mathbf{\mathbf{k}}4}]^{T},$ (15)
and $h(k)=$
$\left(\begin{array}[]{cccc}2J-\frac{J^{\prime}}{2}&\frac{J^{\prime}}{2}e^{\frac{i(k_{x}-k_{y})}{2}}&J\cos(\frac{k_{x}}{2})&J\cos(\frac{k_{y}}{2})\\ \frac{J^{\prime}}{2}e^{\frac{-i(k_{x}-k_{y})}{2}}&2J-\frac{J^{\prime}}{2}&J\cos(\frac{k_{y}}{2})&J\cos(\frac{k_{x}}{2})\\ J\cos(\frac{k_{x}}{2})&J\cos(\frac{k_{y}}{2})&2J-\frac{J^{\prime}}{2}&\frac{J^{\prime}}{2}e^{\frac{-i(k_{x}+k_{y})}{2}}\\ J\cos(\frac{k_{y}}{2})&J\cos(\frac{k_{x}}{2})&\frac{J^{\prime}}{2}e^{\frac{i(k_{x}+k_{y})}{2}}&2J-\frac{J^{\prime}}{2}\end{array}\right).$ (16)
Then we can diagonalize $H_{\text{LSW}}$ via the Bogoliubov transformation with $\Psi(\mathbf{k})=T_{\mathbf{k}}\Phi(\mathbf{k})$}, where
$\Phi(\mathbf{k})=[\beta_{\mathbf{k}1},\beta_{\mathbf{k}2},\beta_{\mathbf{k}3}^{\dagger},\beta_{\mathbf{k}4}^{\dagger}]^{T},$ (17)
is the diagonalized basis and $T_{\mathbf{k}}$ is the transformation matrix.
The general form of spin-wave dispersion is complex but can be easily computed numerically. Here we present the analytical form of the dispersions of the two Goldstone modes along the $\Gamma$-M direction near the $\Gamma$ (and equivalently, M) point. Up to the second-order terms in the wave vector $k_{x}=k_{y}$, the dispersions for $J\neq J^{\prime}$ are
$\begin{matrix}E_{+}&=\sqrt{\frac{2J^{2}(J-J^{\prime})}{2J-J^{\prime}}}k_{x}+\frac{J^{2}J^{\prime}}{2(2J-J^{\prime})^{2}}k_{x}^{2},\\ \\ E_{-}&=\sqrt{\frac{2J^{2}(J-J^{\prime})}{2J-J^{\prime}}}k_{x}-\frac{J^{2}J^{\prime}}{2(2J-J^{\prime})^{2}}k_{x}^{2},\end{matrix}$ (18)
respectively. Note that the coefficients of the linear-in-$k_{x}$ terms are the same, while the splitting caused by the altermagnetism appears in the subleading $k_{x}^{2}$ terms.
At $J=J^{\prime}$ where the AFM ground state becomes unstable in the LSW theory, these two magnon bands are softened with dispersion relations
$\begin{matrix}E_{+}&=\frac{Jk_{x}^{2}}{2},\\ \\ E_{-}&=0.\end{matrix}$ (19)
B.2 Bond operator theory for the triplet excitations in the plaquette phase
In the plaquette phase, the spin rotational symmetry preserves and the LSW theory is not applicable. To derive the low-lying triplet excitations in this phase, we project the Hamiltonian onto the low-energy subspace composed of the singlet ground state and the lowest triplet excitations.
Define a vacuum $|0\rangle$ state and four boson operators which yield the four physical states (one singlet and three triplets) by $|s\rangle=s^{\dagger}|0\rangle$, $|t_{\alpha}\rangle=t_{\alpha}^{\dagger}|0\rangle$, respectively sachdev1990bond ; zhitomirsky1996valence . The projection operators are expressed $Z^{st_{\alpha}}=s^{\dagger}t_{\alpha}$, $Z^{t_{\alpha}t_{\beta}}=t_{\alpha}^{\dagger}t_{\beta}$. The spins represented via these boson operators are
$S_{l}^{\alpha}=\frac{(-1)^{l}}{\sqrt{6}}(s^{\dagger}t_{\alpha}+t_{\alpha}^{\dagger}s)-\frac{i}{4}\sum_{\beta,\gamma}\epsilon_{\alpha\beta\gamma}t_{\beta}^{\dagger}t_{\gamma}$ (20)
where $\alpha,\beta,\gamma=x,y,z$ denote the spin components and $l$ is the sublattice index $(l=1,2,3,4)$. Note that the Hilbert space of the bosons is larger than the original plaquette Hilbert space and includes unphysical states. To restrict the boson Hilbert space to its physical sector, a hard-core constraint must be imposed: $s^{\dagger}s+\sum_{\alpha}t_{\alpha}^{\dagger}t_{\alpha}=1$. By condensing the singlets $s^{\dagger}=s\approx\langle s\rangle$, we rewrite the Hamiltonian in terms of triplet operators
$H=H_{0}+\sum_{\mathbf{k}\in{\text{MBZ}}}\Psi(\mathbf{k}){}^{\dagger}h(\mathbf{k})\Psi(\mathbf{k})+const.,$ (21)
where $H_{0}=E_{s}s^{\dagger}s+\sum_{\alpha}E_{\alpha}t_{\alpha}^{\dagger}t_{\alpha}$ is the energy of a single plaquette and
$\Psi(\mathbf{k})=[t_{\mathbf{k}x},t_{\mathbf{k}y},t_{\mathbf{\mathbf{k}}z},t^{\dagger}_{\mathbf{-k}x},t^{\dagger}_{\mathbf{-k}y},t^{\dagger}_{\mathbf{\mathbf{-k}}z}]^{T}$ (22)
Then following the standard Bogoliubov diagonalization we can derive the dispersion of the lowest triplet as
$E=\sqrt{J^{2}+\frac{2}{3}J(J^{\prime}-2J)(\cos(k_{x})+\cos(k_{y}))}.$ (23)
We compare the triplet dispersion with that calculated by using the tensor network method in Fig. 3 of the main text, and they agree well. The small difference is likely caused by the finite bond dimension $D$ in the tensor network calculation.
Appendix C Numerical results on the spin excitation spectra
Figure 6: Total spin DSF ($\mathcal{S}^{xx+yy+zz}$) in the AFM phase of the SS model calculated by iPEPS method with $D=3$ (in (a) and (b)) and $D=2$ (in (c) and (d)), at $J/J^{\prime}=0.9$ (in (a) and (c)) and $J/J^{\prime}=1.1$ (in (b) and (d)).
Figure 7: Finite $D$ scaling of the spin gap of the Goldstone modes (at M point) in the AFM phase of the SS model at $J/J^{\prime}=1.1$. The dashed line shows a best fit to the gap, which vanishes in the large $D$ limit.
Here we show numerical results on the spin excitation spectra in additional to those in the main text. The spectra illustrate the spin DSF
$\mathcal{S}^{\alpha\beta}(\mathbf{q},\omega)=\frac{1}{N}\sum_{i,j}\int dte^{i\mathbf{q}\cdot(\mathbf{r}_{i}-\mathbf{r}_{j})}e^{i\omega t}\langle S_{i}^{\alpha}(t)S_{j}^{\beta}(0)\rangle$ (24)
with spin components $\alpha,\beta$. In Fig. 2(b) we have shown the spectra in the AFM phase at $J/J^{\prime}=1.1$ calculated in the tensor network approach with bond dimension $D=3$. In Fig. 6, we show calculated spectra for several other $J/J^{\prime}$ values and with different bond dimension $D$. One sees that the spectra show similar features with two split magnon bands, which are the Goldstone modes, at low energy and several longitudinal modes at high energy. Interestingly, the spin gap of the Higgs mode decreases with decreasing $J/J^{\prime}$ toward the PVBS-to-AFM transition, as discussed in the main text.
As an effect of finite $D$, the Goldstone modes exhibit a finite gap. We have calculated the spin gaps with other $D$ values. As shown in Fig. 7, the spin gap reduces with increasing $D$ and eventually vanishes in the large $D$ limit. This indicates that the gapless feature of Goldstone modes is captured by the tensor network calculation in this limit.
Figure 8: (a): Difference of the transverse DSF $\Delta\mathcal{S}=\mathcal{S}^{+-}(\mathbf{q},\omega)-\mathcal{S}^{-+}(\mathbf{q},\omega)$ in the AFM phase of the SS model at $J/J^{\prime}=1.1$ in the LSW theory. (b): Same as (a), but calculated by the tensor network method. In either case, the opposite sign of $\Delta\mathcal{S}$ reflects the opposite chirality of the magnon modes.
To show that the chiral magnon excitations in the AFM phase of the SS model, we calculate the difference of the transverse component of the DSF $\Delta\mathcal{S}=\mathcal{S}^{+-}(\mathbf{q},\omega)-\mathcal{S}^{-+}(\mathbf{q},\omega)$. Note that $S_{i}^{+}S_{j}^{-}-S_{i}^{-}S_{j}^{+}\propto(\mathbf{S}_{i}\times\mathbf{S}_{j})_{z}$, so that the sign of $\Delta\mathcal{S}$ is able to detect the chirality of the spin excitations. In Fig. 8 we compare results of $\Delta\mathcal{S}$ calculated by LSW theory and tensor network method. Both results show a finite splitting between the two Goldstone modes with opposite sign of $\Delta\mathcal{S}$. Also note that the magnon bands along $\Gamma$-M and $\Gamma$-M also have opposite sign. Interestingly, we find that the optical magnon bands also exhibit opposite chirality. These results verify the existence of chiral magnons in the AFM state, which is caused by the symmetry of altermagnetism.
Figure 9: (a): Density of states (DoS) with excitation energy at several $J/J^{\prime}$ values in the AFM phase. (b): Same as (a) but in the plaquette phase.
Besides the collective magnon excitations and the longitudinal modes, the tensor calculation also resolves continuum excitations. The spectral weights of these excitations are much smaller than the collective modes but as shown in Fig. 9, they cause a broad peak in the density of states (DoS) at high energy. In the AFM phase, this broad peak appears in energies above the Higgs gap (above $\sim 2J^{\prime}$ at $J/J^{\prime}=1.1$), and the peak energy decreases with varying $J/J^{\prime}$ toward the PVBS-to-AFM transition. This behavior is consistent with the emergence of deconfined excitations although it is hard to tell whether the continuum has a spinon or magnon origin. In the plaquette phase, we can also identify continuous excitations at energies above the triplet gap (above $\sim J^{\prime}$ at $J/J^{\prime}=0.7$) in the DoS plots (Fig. 9(b)). The behavior of energies of these excitations is similar to that in the AFM phase.
Figure 10: (a): Field dependence of gaps of several low-energy excitations in the plaquette phase of the SS model, calculated by the tensor network method. The multiplets split into $S^{z}=+1,-1,0$ states. (b): Field dependence of low-energy excitations of a 4-spin cluster with nearest neighbor AFM interactions, calculated by the ED. The splitting indicates that the 7-fold multiplet contains two degenerate triplets and one singlet.
To understand the singlet and the 7-fold multiplet modes above the lowest triplet mode in the plaquette phase resolved by the tensor network calculation in the main text, we apply a small magnetic field to split the multiplets. We also compare our tensor network results with those from exact diagonalization on a four-spin plaquette with nearest neighbor Heisenberg interaction. As shown in Fig. 10(a), the 7-fold quasi-degenerate excitations splits under a small field. The gaps of three modes are unchanged under the field, while the gaps of two modes increase with the field, gaps of another two modes decrease with the field. Note that the slops of the gaps are the same for these modes, and they are also same to the slops of the split lowest triplet excitations. These indicate that the 7-fold multiplet actually consists of one singlet and two degenerate triplets. This conclusion is further confirmed by our exact diagonalization calculation shown in Fig. 10(b). At zero field, there are indeed 7-fold degenerate multiplet excitations above the triplet ones. With applying a small magnetic field, The 7-fold multiplet splits into a pair of triplets and one singlet. Note that the singlet excitation in between the lowest triplet and the 7-fold multiplet seen in the tensor network result in Fig. 3 of the main text does not show up in our 4-spin cluster exact diagonalization calculation. It is then interpreted as an inter-plaquette bound state.
Figure 11: Evolution of gaps at the M point of the triplet (T) and singlet (S) modes in the PVBS phase, as well as the Higgs (H), magnon (M), and singlet (S) modes in the AFM phase, with $J/J^{\prime}$ near the PVBS-to-AFM transition in the iPEPS calculation with $D=3$ (in (a)) and $D=4$ (in (b)). In either case we observe softening of all collective modes when approaching the transition point.
In Fig. 4(a) of the main text we discuss the evolution of various collective excitations with varying $J/J^{\prime}$. These include triplet and singlet modes in the PVBS phase, and the Higgs, magnon, and singlet modes in the AFM phase. It is found in Fig. 4(c) that the gaps of triplet and singlet modes in the PVBS phase, together with those of the Higgs and singlet modes in the AFM phase, drop rapidly to small values close to zero, and become quasi-degenerate with the magnon (Goldstone) modes when approaching the PVBS-to-AFM transition. Here in Fig. 11 we show results with smaller bond dimensions $D=3$ and $D=4$. We observe the same behavior that all these collective modes are softened near the transition. Note that the gap between the Higgs and magnon modes is still sizable at the transition point for $D=3$ and $D=4$. But as shown in Fig. 4(c) and (e), this gap drops rapidly with increasing $D$ and eventually closes in the large $D$ limit. This behavior strongly suggests the emergence of a DQCP at the PVBS-to-AFM transition.